Please check to make sure each of the following reactions is correct (with correct charges and coefficients!). Also, please identify any amphoteric ions in the equations.

CuSO4(aq) + 2HCl(aq) =CuCl2(aq) + H2SO4(aq)

CuCl2(aq) + H2O(aq) = [Cu(H2O)2Cl2](aq)

CoCl2(aq) + 2HCl(aq) =CoCl42-(aq) + H2(g)

CoCl42-(aq) + 6H2O =Co(H2O)62++ 4Cl-(aq)

Ni(NO2)2(aq) + 2H2O(aq) =Ni(OH)2(aq) + 2HNO2(aq)

CuSO4(aq) + 2NaOH(aq) =Na2SO4(aq) + Cu(OH)2(s)

KSCN(aq) + CuSO4(aq) =KSO4(aq) + CuSCN+(aq)

CuSCN+(aq)+ 2NaOH(aq) =Cu(OH)2(s)+ 2Na+(aq) + SCN-(aq)

CuSO4(aq) + Na2C4H4O6(aq) =CuC4H4O6(s) + Na2SO4(aq)

NaOH(aq) + CuSO4(aq) =Cu(OH)2(s) + Na2SO4(aq)

NiCl2(aq) + 6 NH4OH(aq) =6 H2O(aq)  + [Ni(NH3)6]Cl2(aq)

2NaOH(aq) + NiCl2(aq) + 2KSCN(aq)=2NaSCN(aq) + Ni(OH)2(s) + 2KCl(aq)

NiCl2(aq) + Na2C2O4(aq) =NiC2O4(aq) + 2NaCl(s)

NiCl2(aq) + 2NaOH(aq) = 2NaCl(s) + Ni(OH)2(aq)

CoCl2(aq) + 2NH4OH(aq) =Co(OH)2(aq) + 2NH4Cl(s)

Co(OH)2(aq) + 2NaOH(aq) =Na2[Co(OH)4](s)

CoCl2(aq) + KSCN(aq) = CoSCN(aq)  + KCl2(aq)

KCl2(aq) + NaOH(aq) = NaCl2(s) + KOH(aq)

CoCl2(aq) + 2NaOH(aq) = Co(OH)2(aq) + 2NaCl(aq)

CoCl2(aq) + Na2C4H4O6(aq)= CoC4H4O6(aq) + 2NaCl(aq)

CoC4H4O6(aq) + 2NaOH(aq) =Co(OH)2(s) + Na2C4H4O6(aq)
Answer #1

The following reactions are corrected:

2KSCN (aq) + CuSO4 (aq) =K2SO4 (aq) + Cu(SCN)2 (aq)

2 NaOH (aq) + CuSO4 (aq) =Cu(OH)2 (s) + Na2SO4 (aq)

CoCl2 (aq) + 2 KSCN (aq) = Co(SCN)2 (aq)  + 2 KCl (aq)

KCl (aq) + NaOH (aq) = NaCl (s) + KOH (aq)

Amphoteric ions:

Ions behaving as both acids and bases. Eg: HCO3-, HSO4- etc.

Here in these equations no amphoteric ions are present.